In an elevator that conveniently has a scale, a lady of 65.0 kg mass, was happy to read 442 N for her weight on the scale. Determine

  1. (a) the direction of the elevator’s acceleration.

    (b) the magnitude of the elevator’s acceleration
Answer #1

here,

the mass of lady , m = 65 kg

the weight of lady , W = m * g

W = 65 * 9.81 N = 637.65 N

a)

as the apparent weight of lady is less than the weight of lady

the direction of acceleration of elevator must be Downwards

b)

let the magnitude of acceleration be a

the apparent weight , W' = 442 N = m * ( g - a)

442 = 65 * ( 9.81 - a)

solving for a

a = 3.01 m/s^2

the magnitude of acceleration is 3.01 m/s^2